# Refusal Tokens: A Simple Way to Calibrate Refusals in Large Language Models

### v1

@@ 7 Discussion
Another aspect of refusals to consider is adversarial attacks. Although we assume that the user in these settings is not acting maliciously, an individual may optimize the refusal tokens directly optimize on short strings like “Sure here’s,..” such as Shin et al. (2020); Wen et al. (2023); Zou et al. (2023); Zhu et al. (2024). However, these attacks are well-studied in the community (Alon & Kamfonas, 2023; Jain et al., 2023; Zhou et al., 2024). A more specific threat model involves scenarios where a user places the respond token either at the end of the input or the beginning of a response. In an API setting, such attacks can be filtered out.
Table 4: The counts of response tokens or refusal tokens generated and what the model generation was labeled. Left shows the counts for a single refusal token under default sampling parameters. Right shows the counts for category refusal tokens under default sampling parameters.
Response
Label
Refusal Token
Generated
Response Token
Generated
Refused 1019 46
Responded 29 277
Response
Label
Refuse Cat.
Generated
Respond Token
Generated
Refused 945 68
Responded 43 315
The ability of a model to refuse queries–whether due to toxicity, limitations, or other reasons–is crucial for developing safer and more trustworthy LLMs. To advance this, we need to understand how and why models generalize across different contexts, which requires the appropriate data. While some datasets, such as Brahman et al. (2024), provide broad coverage, there remains a gap in preference data and multi-turn evaluations, complicating the task of generalizing single-turn results to multi-turn interactions. Thus, we need additional data to better understand this property of LLMs.
Nevertheless, adding a refusal token during fine-tuning offers several benefits. When the model generates the token, it associates a softmax probability of refusal with the query. At test-time, the refusal token allows for adjusting the refusal rate. Moreover, by applying the refusal token to specific categories, the distribution can be controlled, and thresholding techniques can further improve the F1 scores of refusal rates. Additionally, these tokens can be modified in various ways during testing, such as using logit bias, category-specific thresholding, or sum thresholding, highlighting their flexibility. Therefore, without retraining language models, refusal tokens offer the advantage of test-time control, benefiting both users and API providers.
@@ Appendix A Appendix
A.1 XSTest
@@
Since the original CoCoNot and temporal setting in the test set is reflective of the train set (as they come from the same source), we suspect the behavior of the token is better when the train and test distributions are more aligned in terms of wording. In this new setting, we train on the CoCoNot train set and evaluate on the XSTest test set. Thus, we want to confirm that some of the capabilities such as turning off the token to reduce overall refusal rates and the out-of-the-box benefits are present. From Table 5, we see that adding the refusal tokens improves the full refusal rate on unsafe and lowers the safe refusal rate by 1% in either direction. Additionally, adding the category refusal tokens decreases the safe refusal rate by over 5% and slightly reduces the refusal rate on the unsafe questions by about 0.5%. When analyzing the outputs for the difference in 5% for category tokens versus refusal tokens, we observed that different category tokens were utilized providing a non-safety reason that yielded in gpt4 marking them as a compliant response. Additionally, to confirm that the tokens can affect refusal rates for this set of prompts, we experiment with only producing the respond token, or turning off the refusal tokens. We find that this token reduces the overall refusal rate by up about 5% for model that contain category tokens and about 10% for the model trained with a single refusal token. This result further validates the token’s ability to control refusal rates at test-time.
@@
Setting Refusal Rate on Safe Prompts Refusal Rate on Unsafe Prompts
Baseline 17.2% (4.4%) 89.0% (0.00%)
+ Refusal Tokens 16.4% (4.4%) 90.5% (0.00%)
+ Refusal Tokens OFF 5.6% (3.2%) 63.5% (0.00%)
+ Category Tokens 12.0% (1.6%) 88.5% (0.00%)
+ Category Tokens OFF 6.8% (1.2%) 72.5% (0.00%)
A.2 Additional Experiments For Out-of-the-Box Training
@@
A.3 Temporal Training Data and TempEval
@@
System Prompt === “{task_prompt}. The passage is from the year {year}. Be specific and ONLY return the question.” User Prompt === “{article}”
@@ A.4 Prompt For CoCoNot
A.5 Additional Training Details.
We used the codebase from Tunstall et al. (2023) and the hyperparameters as well. We trained the models with bfloat16, Flash Attention-2 (Dao, 2024), and packing. We used a learning rate of ${2.0e-5}$ with cosine decay. Additionally, hyperparameter details can be found in Tunstall et al. (2023) at https://github.com/huggingface/alignment-handbook. We altered the sequence length for training from $2048$ to $1024$. For Alpaca, we trained for three epochs and one epoch for UltraChat. We used the chat template from llama-3 Instruct. Additionally, we the chat template from llama-3. The majority of training runs were completed on $8$ Nvidia A100 $40$GB.
A.6 Thresholding Algorithms

### v2

@@ 9 Ethics Statement
Adding the refusal token to the model may allow for the adversarial attacks that try to jailbreak the model such as those suggested by Shin et al. (2020); Wen et al. (2023); Zou et al. (2023); Zhu et al. (2024) where an adversary optimizes an adversarial prompt on a short string to jailbreak the model. Although we assume that the user in these settings is not acting maliciously, an individual may optimize the refusal tokens directly optimize on short strings like “Sure here’s,..” such as Shin et al. (2020); Wen et al. (2023); Zou et al. (2023); Zhu et al. (2024). However, these attacks are well-studied in the community (Alon & Kamfonas, 2023; Jain et al., 2023; Zhou et al., 2024). A more specific attack to the refusal token involves scenarios where a user places the respond token either at the end of the input or the beginning of a response. In an API setting, the user should not have access to such a token.
10 Acknowledgements
Neel Jain and Ashwinee Panda were supported by Captial One Bank during the majority of work. Additionally, this work was made possible by the ONR MURI program, DAPRA TIAMAT, the National Science Foundation (IIS-2212182), and the NSF TRAILS Institute (2229885). Commercial support was provided by Capital One Bank and Open Philanthropy.
@@ Appendix A Appendix
A.1 Sum Thresholding
Improving F1 scores with sum thresholding. The sum thresholding scheme can be considered where controlling individual categories is not of interest. Particularly, we sweep the thresholds of a model trained with UltraChat, CoCoNot refusal messages, and CoCoNot contrast training examples. In Figure 6, by sweeping the thresholds between 0 and $1$ in increments of $0.1$, a threshold of $0.6$ yields the best F1 score over sampling. This experiment further shows that category tokens can be manipulated at test-time by either thresholding, individual or sum, scheme for better F1 performance or different needs. Overall, using multiple tokens provides greater flexibility and steerability for the user than a single refusal token. However, if a user does not require this level of flexibility or prefers not to add many new tokens to the vocabulary, a single refusal token remains an excellent solution compared to adding individual tokens for different category tokens for controlling the model’s refusal rate, as shown in Figure 2. Ultimately, the choice depends on the user’s specific preferences and requirements.
Figure 6: Sum thresholding is another way to effectively utilize the category tokens at test-time. (Upper left) F1 scores on CoCoNot evaluation, (upper right) average of the refusal rates for refusal categories in the CoCoNot evaluation, and (bottom) the refusal rate the contrast category in the CoCoNot evaluation as the threshold is swept. The refusal token is emitted if the sum of the scores for all category tokens exceeds the threshold. At a threshold of $T=0.6$, the F1 Score is highest at $0.946$ up from $0.938$, cutting the error rate by $\sim 12\%$.
A.2 XSTest
@@
Since the original CoCoNot and temporal setting in the test set is reflective of the train set (as they come from the same source), we suspect the behavior of the token is better when the train and test distributions are more aligned in terms of wording. In this new setting, we train on the CoCoNot train set and evaluate on the XSTest test set. Thus, we want to confirm that some of the capabilities such as turning off the token to reduce overall refusal rates and the out-of-the-box benefits are present. From Table 5, we see that adding the refusal tokens improves the full refusal rate on unsafe and lowers the safe refusal rate by 1% in either direction. Additionally, adding the category refusal tokens decreases the safe refusal rate by over 5% and slightly reduces the refusal rate on the unsafe questions by about 0.5%. When analyzing the outputs for the difference in 5% for category tokens versus refusal tokens, we observed that different category tokens were utilized providing a non-safety reason that yielded in GPT-4 marking them as a compliant response. Additionally, to confirm that the tokens can affect refusal rates for this set of prompts, we experiment with only producing the respond token, or turning off the refusal tokens. We find that this token reduces the overall refusal rate by up about 5% for model that contain category tokens and about 10% for the model trained with a single refusal token. This result further validates the token’s ability to control refusal rates at test-time.
@@
Setting
Refusal Rate on
Safe Prompts
Refusal Rate on
Unsafe Prompts
Baseline 17.217.2% (4.44.4%) 89.089.0% (0.000.00%)
+ Refusal Tokens 16.416.4% (4.44.4%) 90.590.5% (0.000.00%)
+ Refusal Tokens OFF 5.65.6% (3.23.2%) 63.563.5% (0.000.00%)
+ Category Tokens 12.012.0% (1.61.6%) 88.588.5% (0.000.00%)
+ Category Tokens OFF 6.86.8% (1.21.2%) 72.572.5% (0.000.00%)
A.3 Additional Experiments For Out-of-the-Box Training
@@
A.4 Temporal Training Data and TempEval
@@
System Prompt == “{task_prompt}. The passage is from the year {year}. Be specific and ONLY return the question.” User Prompt == “{article}”
@@ A.5 Prompt For CoCoNot
A.6 Additional Training Details.
We used the codebase from Tunstall et al. (2023) and the hyperparameters as well. We trained the models with bfloat16, Flash Attention-2 (Dao, 2024), and packing. We used a learning rate of 2.0e5{2.0e-5} with cosine decay. Additionally, hyperparameter details can be found in Tunstall et al. (2023) at https://github.com/huggingface/alignment-handbook. We altered the sequence length for training from $2048$ to $1024$. For Alpaca, we trained for three epochs and one epoch for UltraChat. We used the chat template from llama-3 Instruct. Additionally, we the chat template from llama-3. The majority of training runs were completed on $8$ Nvidia A100 4040GB.
A.7 Thresholding Algorithms